OpenStudy (anonymous):

sketch the region enclosed by the given curves and find its area: y=12-x^2 y=x^2-6

OpenStudy (stamp):

Quadratic equations, so we are dealing with parabolas. Find the extrema of each vertex, and identify it's concavity. Then you can begin to visualize each function and begin to plot. Once your plot is graphed we can consider finding the area.

OpenStudy (anonymous):

Would it be the integral from -3 to 3 of (12-x^2) - (x^2-6)?

OpenStudy (stamp):

\[f_{1}(x)=12-x^2\]\[f_{2}(x)=x^2-6\] @alma1apple I don't know, I don't start solving the integral for the area until I have graphed my functions.